# Geometry and Optimization of Shallow Polynomial Networks

### v1

@@ 3.2 Critical points of the distance function
When $X$ is (the smooth locus of) a nonlinear algebraic variety, the complex Zariski closure of the extended focal locus $F_{X}$ is a hypersurface [10, Theorem 1]. In particular, there exists a single polynomial equation $P(t,\Sigma)=0$ in both the teacher and inner product that must be satisfied when there is a transition in the qualitative properties of the landscape (equivalently, when this equation is not satisfied, the qualitative landscape locally does not change). Moreover, if the inner product $\Sigma=\Sigma_{\mathcal{D}}$ is induced by a distribution $\mathcal{D}$ as above, then we also obtain a polynomial $Q(t,{\bm{\mu}}):=P(t,\Sigma_{\mathcal{D}})$ in the teacher and moments $\bm{\mu}$ of $\mathcal{D}$. In this setting, the equation $Q(t,{\bm{\mu}})=0$ precisely defines all “unstable data” for the optimization landscape. We refer to $P$ and $Q$ as the teacher-metric discriminant and the teacher-data discriminant, respectively. We also note that Theorem 25 can in some cases be applied to nonsmooth function spaces stratum-by-stratum, as we do in Section 4 for determinantal varieties.
@@
Finally, the following result from [37] shows that the signature of the Hessian of the distance function has an intuitive geometric description in terms of the focal locus. We will use this fact to easily compute the signature of the Hessian of the loss function for quadratic networks.
@@
For any $t\in\mathbb{R}^{n}\setminus F_{X,\Sigma}$ and $s\in{\rm Crit}(L_{t,\Sigma}|_{X})$, the index of $L_{t,\Sigma}|_{X}$ at $s$ (i.e., number of negative eigenvalues of the Hessian $H_{L_{t,\Sigma}}$) is equal to the number of focal points of $(X,s)$ which lie in the segment $s$ from $t$, each counted with multiplicity.
@@
In the proof of Proposition 23 we noted that $\frac{\partial^{2}L}{\partial u_{i}\partial u_{j}}=2\left(\frac{\partial x}{\partial u_{i}}\cdot\frac{\partial x}{\partial u_{i}}-\frac{\partial^{2}x}{\partial u_{i}\partial u_{j}}\cdot(t-s)\right)$, where $x(u)$ is a local parameterization of $X$. Without loss of generality we can assume that $\frac{\partial x}{\partial u_{i}}\cdot\frac{\partial x}{\partial u_{i}}$ is the identity. If $t-s=\gamma w$, with $\gamma\in\mathbb{R}$ and $\|w\|=1$, then the index of the Hessian is the number of eigenvalues $\frac{\partial^{2}x}{\partial u_{i}\partial u_{j}}\cdot w$ (or principal curvatures, cf. Remark 24) which are greater than $1/\gamma$. On the other hand, $t$ is a focal point if and only if $1/\gamma$ is an eigenvalue, and its multiplicity is equal to the multiplicity of $1/\gamma$ as an eigenvalue. ∎
@@ 3.3 Examples
In this setting, the teacher-metric discriminant is a homogeneous polynomial $P(t_{00},t_{01},t_{11},m_{0000},m_{0001},m_{0011},m_{0101},m_{0111},m_{1111})$ of bidegree $(6,12)$ with $5938$ terms. The Frobenius inner product corresponds to $m_{0000}=m_{1111}=1$, $m_{0101}=\frac{1}{2}$, and $m_{0001}=m_{0011}=m_{0111}=0$. Under these conditions, the discriminant specializes to
@@
which is (a power of) the discriminant of the characteristic polynomial of $T$, as noted in [16]. For distribution-induced inner products, $m_{ijkl}=\mathbb{E}[x_{i}x_{j}x_{k}x_{l}]$, and in particular $m_{0011}=m_{0101}$. Imposing this condition on $P$ reduces it to a polynomial with $2060$ terms. If we further assume a symmetric i.i.d. distribution and define $\mu_{4}=m_{0000}=\mathbb{E}[x^{4}]$ and $\mu_{2}^{2}=m_{0011}=m_{0111}=\mathbb{E}[x^{2}]^{2}$, the full discriminant is given by
$+\mu_{4}^{12}t_{00}^{3}t_{11}^{3}$
Similar to the previous example, slices of the teacher-metric discriminant can be visualized as varying focal loci (Figure 2, right).
@@ 4 Networks with Quadratic Activations
We now focus on polynomial networks $f_{\mathcal{W}}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ with quadratic activations $\rho(z)=z^{2}$. For $\mathcal{W}=(\alpha,W)\in\mathbb{R}^{r}\times\mathbb{R}^{r\times n}$, we write
@@ 4.1 Critical points of smooth loss functions
We begin by specializing and refining some of the general results on the critical points of smooth loss functions from Section 2 to the case of $d=2$. In the following, we write $\mathcal{S}(r;n)$ for the set of $n\times n$ symmetric matrices with rank exactly $r$. The set $\mathcal{S}(r;n)$ is a smooth manifold of dimension $\frac{1}{2}r(2n-r+1)$ and the embedded tangent space at a point $S\in\mathcal{S}(r;n)$ is $T_{S}\mathcal{S}(r;n)=\{SX+X^{\top}S\colon X\in\mathbb{R}^{n\times n}\}$ [24].
@@
We show that $T_{S}\mathcal{S}(k;d)^{\perp}=\{A\in S^{2}(\mathbb{R}^{d})\colon A\cdot S=0\}$. Indeed, if $A\cdot S=0$, then from the cycle property of the trace operator $\langle A,SX+XS\rangle_{F}=tr(ASX)+tr(AX^{\top}S)=0$ for any $X\in\mathbb{R}^{d\times d}$. Conversely, if $\langle A,SX+X^{\top}S\rangle_{F}=0$ for all $X$, then $As_{i}=0$ for all column $s_{i}$ of $S$ which means that $A\cdot S=0$. ∎
@@
If If $\ell:{\rm Sym}^{2}(\mathbb{R}^{n})\rightarrow\mathbb{R}$ is any smooth function and $L=\ell\circ\tau_{r}$, then for any $\mathcal{W}=(\alpha,W)\in\mathbb{R}^{r}\times\mathbb{R}^{r\times n}$ we have that
@@
If $L=\ell\circ\tau_{r}$ and $\mathcal{W}=(\alpha,W)$ is a critical point for $L$ such that $S=\tau_{r}(\mathcal{W})$, then $\nabla\ell(S)\cdot S=0$ holds. In particular, if $r^{\prime}=rk(S)$, we have that $S$ is a critical point of the restriction of $\ell$ to $\mathcal{S}(r;n)$. More generally, the image of the entire critical set $\tau_{r}(Crit(L))$ is equal to $\bigcup_{r^{\prime}\leq r}Crit(\ell|_{\mathcal{S}(r^{\prime};n)})$.
@@
The first statement is a consequence of the first part of Proposition 30, noting that if $\mathcal{W}$ is a critical point for $L$ then $0=W\cdot\nabla_{W}L(\mathcal{W})=4\cdot S\cdot\nabla L(S)$. The second statement then follows from Proposition 29. For the last statement, consider $S$ such that $rk(S)=r^{\prime}<r$ and $\nabla\ell(S)\cdot S=0$, so $S\in Crit(\ell|_{\mathcal{S}(r^{\prime};n)})$. A parameterization $S=\tau_{r}(\mathcal{W})$ such that $\alpha_{i}\neq 0$ for exactly $r^{\prime}$ indices will be such that then $\nabla L(\mathcal{W})=0$ (again from Proposition 30). It enough to note that any $S$ with $rk(S)=r^{\prime}<r$ admits a representation in parameter space with exactly $r^{\prime}$ non-zero neurons. ∎
@@ 4.2 Teacher-student problems
The expression (28) essentially appears in [21, Theorem 3.1(b)], but as a norm rather than as an inner product. Using (27) and the assumptions on $\mathcal{D}$, we have that
@@
From Corollary 31, it is sufficient to describe critical points of the loss restricted to the low-rank manifold $\mathcal{S}(r;n)$. Thus, we consider teacher-student problem in function space:
@@
[…]
For the Frobenius inner product, the optimization landscape is described by the following version of Eckart-Young’s Theorem, shown in [24]. We present here a much shorter geometric proof using the general description of critical points for distance functions given in Section 3 (the proof in [24] consists in more than 3 pages of computations). We make use of the following intuitive description of the focal locus.
@@
Let $X=\mathcal{S}(r,n)$ and let $\Sigma$ correspond to the Frobenius or Gaussian norm. Given $S\in X$, consider $A\in{\rm Sym}^{2}(\mathbb{R}^{n})$ such that $\{T_{\gamma}=S+\gamma A\colon\gamma\geq 0\}$ is a $\Sigma$-orthogonal ray to $X$. Assume further that there exists a matrix along this ray with distinct eigenvalues. Then $T_{\gamma}$ is a focal point for $(X,S)$ if and only if it has a repeated eigenvalue. More precisely, the multiplicity of $T_{\gamma}=\sum_{i=1}^{n}\sigma_{i}(u_{i}\otimes u_{i})$ as a focal point is given by #$\#\{(j,k)\colon j<k,\sigma_{j}=\sigma_{k}\}$.
@@
  • The function $h_{T}|_{\mathcal{S}(r;n)}:\mathcal{S}(r;n)\rightarrow\mathbb{R}$ has exactly $\binom{n}{r}$ critical points, given by $S_{\mathcal{I}}=U\Sigma_{\mathcal{I}}U^{T}$ for any $\mathcal{I}\subset\{1,\ldots,n\}$ such that #$\#\mathcal{I}=r$.
@@
We have that $S\in{\mathcal{S}(r;n)}$ is a critical point of $h_{T}|_{\mathcal{S}(r;n)}$ if and only if $\langle T-S,XS+SX\rangle=0$ for all $X\in{\rm Sym}^{2}(\mathbb{R}^{n})$, which in turn is equivalent to $S\cdot(T-S)=0$ by Proposition 29. Now let $S=\sum_{i=1}^{r}\sigma_{i}^{\prime}(u_{i}^{\prime}\otimes u_{i}^{\prime})$ and $(T-S)=\sum_{j=1}^{s}\sigma_{j}^{\prime\prime}(u_{j}^{\prime\prime}\otimes u_{j}^{\prime\prime})$ be eigendecompositions of $S$ and $T-S$ with $\sigma_{i}^{\prime}\neq 0$ and $\sigma_{j}^{\prime\prime}\neq 0$. The row/column spaces of $S$ and $T-S$ are spanned by the vectors $u_{i}^{\prime}$ and $u_{j}^{\prime\prime}$, respectively. From the condition $S(T-S)=0$, we deduce that $S$ is a critical point if and only if $u_{i}^{\prime}\cdot u_{j}^{\prime\prime}=0$ for all $i,j$ in the two decompositions, which means that
@@
is an eigendecomposition of $T$. This shows that critical points are of the form $S=U\Sigma_{\mathcal{I}}U^{T}$ where $U\Sigma U^{T}$ is an eigendecomposition of $T$ and $\mathcal{I}\subset\{1,\ldots,n\}$ is such that #$\#\mathcal{I}=r$.
@@
  • The function $h_{T}|_{\mathcal{S}(r;n)}:\mathcal{S}(r;n)\rightarrow\mathbb{R}$ has exactly $\binom{n}{r}$ critical points, given by $S_{\mathcal{I}}=U\Lambda_{\mathcal{I}}U^{T}$ for any $\mathcal{I}\subset\{1,\ldots,n\}$ such that #$\#\mathcal{I}=r$, with
@@
Assume that $S=\sum_{i=1}^{r}\sigma_{i}^{\prime}(u_{i}^{\prime}\otimes u_{i}^{\prime})$ and $(S-T)=\sum_{i=1}^{n}\sigma_{i}^{\prime\prime}(u_{i}^{\prime\prime}\otimes u_{i}^{\prime\prime})$. Setting $\tau=tr(S-T)$, we have $\nabla h_{T}(S)=\sum_{i=1}^{n}(2\sigma_{i}^{\prime\prime}+\tau)(u_{i}^{\prime\prime}\otimes u_{i}^{\prime\prime})$. From the condition $S\cdot\nabla h_{T}=0$, we deduce that $S$ and $\nabla h_{T}$ have orthogonal eigenvectors, i.e., $u_{i}^{\prime}\cdot u_{j}^{\prime\prime}=0$ holds for all $i$ and all $j$ such that $2\sigma^{\prime\prime}_{j}+\tau\neq 0$. This implies that a decomposition of $(S-T)$ is such that, up to permutation of indices, $u_{i}^{\prime\prime}=u_{i}^{\prime}$ and $\sigma_{i}^{\prime\prime}=-\tau/2$ holds for $i=1,\ldots,r$, while $\sum_{i=r+1}^{n}\sigma^{\prime\prime}_{i}=\tau(r+2)/2$ (so that $tr(S-T)=\tau$). We thus have that an eigendecomposition of $T$ is given by
@@
To prove the second point, we use again Theorem 26 and Proposition 33. The eigenvalues of $M_{\alpha}=\alpha T+(1-\alpha)S$ are the diagonal elements of $\Sigma_{\mathcal{I}}+\alpha\Sigma_{\mathcal{I}^{c}}+(1-\alpha)c_{\mathcal{I}}I_{\mathcal{I}}$ where $c_{\mathcal{I}}=\frac{tr(\Sigma_{\mathcal{I}^{c}})}{r+2}$ and $I_{\mathcal{I}}$ is the diagonal matrix with $I_{\mathcal{I},ii}=1$ if $i\in\mathcal{I}$ and $I_{\mathcal{I},ii}=0$ otherwise. Two eigenvalues of $M_{\alpha}$ are the same if and only there exist $i\in\mathcal{I}$ and $j\in\mathcal{I}^{c}$ such that $\sigma_{i}+(1-\alpha)c_{\mathcal{I}}=\alpha\sigma_{j}$, or equivalently $\alpha=\frac{\alpha_{i}+c_{\mathcal{I}}}{\alpha_{j}+c_{\mathcal{I}}}$. The number of intersections with the discriminant locus is thus equal to the number of pairs $(\sigma_{i},\sigma_{j})\in\mathcal{I}\times\mathcal{I}^{c}$ such that ${(\sigma_{i}+c_{\mathcal{I}})}/{(\sigma_{j}+c_{\mathcal{I}})}\in[0,1]$. ∎
Finally, we prove that using metrics other than Gaussian and Frobenius, there may be exponentially more critical points.
@@
has $(3^{n}-1)/2$ many critical points, whenever $T=\operatorname{diag}(t)$ for $t\in\mathcal{T}$.
Theorem 36 is proven in the appendix. Note that it stands in contrast to the Gaussian and Frobenius cases for low-rank approximation. There, the corresponding optimization problem has merely $n$ critical points generically. The number of critical points $(3^{n}-1)/2$ in Theorem 36 coincides with the “generic ED degree” reported in [16, Example 5.6].

### v2

@@ 3.2 Critical points of the distance function
This result provides a tool to study the qualitative behavior of the optimization landscape as the data varies. Indeed, when $X$ is (the smooth locus of) a nonlinear algebraic variety, the complex Zariski closure of the extended focal locus $F_{X}$ is a hypersurface [catanese2000focal, Theorem 1]. In particular, there exists a single polynomial equation $P(t,\Sigma)=0$ in both the teacher and inner product that must be satisfied when there is a transition in the qualitative properties of the landscape (equivalently, when this equation is not satisfied, the qualitative landscape remains locally unchanged). Moreover, if the inner product $\Sigma=\Sigma_{\mathcal{D}}$ is induced by a distribution $\mathcal{D}$ as above, then we also obtain a polynomial $Q(t,{\bm{\mu}}):=P(t,\Sigma_{\mathcal{D}})$ in the teacher and moments $\bm{\mu}$ of $\mathcal{D}$. The equation $Q(t,{\bm{\mu}})=0$ then precisely defines all “unstable data” for the optimization landscape. We refer to $P$ and $Q$ as the teacher-metric discriminant and the teacher-data discriminant, respectively. More broadly, both are instances of what we call a data discriminant: loci in data spaces that reflect qualitative changes in the optimization of the training loss. We also note that Theorem 26 can in some cases be applied to nonsmooth function spaces stratum-by-stratum, as we do in Section 4 for determinantal varieties which are stratified by matrix rank.
@@
The next result from [milnor2016morse] shows that the signature of the Hessian of the distance function has an intuitive geometric description in terms of the focal locus. We will use this fact to easily compute the signature of the Hessian of the loss function for quadratic networks in Section 4.
@@
For any $t\in\mathbb{R}^{n}\setminus F_{X,\Sigma}$ and $s\in{\rm Crit}(L_{t,\Sigma}|_{X})$, the index of $L_{t,\Sigma}|_{X}$ at $s$ (i.e., number of negative eigenvalues of the Hessian $H_{L_{t,\Sigma}}$) is equal to the number of focal points of $(X,s)$ which lie in the segment from $s$ to $t$, each counted with multiplicity.
@@
In the proof of Proposition 24 we noted that $\frac{\partial^{2}L}{\partial u_{i}\partial u_{j}}=2\left(\frac{\partial x}{\partial u_{i}}\cdot\frac{\partial x}{\partial u_{{j}}}-\frac{\partial^{2}x}{\partial u_{i}\partial u_{j}}\cdot(t-s)\right)$, where $x(u)$ is a local parameterization of $X$. Without loss of generality we can assume that $\frac{\partial x}{\partial u_{i}}\cdot\frac{\partial x}{\partial u_{{j}}}$ is the identity. If $t-s=\gamma w$, with $\gamma\in\mathbb{R}$ and $\|w\|=1$, then the index of the Hessian is the number of eigenvalues $\frac{\partial^{2}x}{\partial u_{i}\partial u_{j}}\cdot w$ (or principal curvatures, cf. Remark 25) which are greater than $1/\gamma$. On the other hand, $t$ is a focal point if and only if $1/\gamma$ is an eigenvalue, and its multiplicity is equal to the multiplicity of $1/\gamma$ as an eigenvalue. ∎
@@ 3.3 Examples
In this setting, the teacher-metric discriminant is a homogeneous polynomial
$P(t_{00},t_{01},t_{11},m_{0000},m_{0001},m_{0011},m_{0101},m_{0111},m_{1111})$
of bidegree $(6,12)$ with $5938$ terms. The Frobenius inner product corresponds to $m_{0000}=m_{1111}=1$, $m_{0101}=\frac{1}{2}$, and $m_{0001}=m_{0011}=m_{0111}=0$. Under these conditions, the discriminant specializes to
@@
which is (a power of) the discriminant of the characteristic polynomial of $T$, as noted in [draisma2013]. For distribution-induced inner products, $m_{ijkl}=\mathbb{E}[x_{i}x_{j}x_{k}x_{l}]$, and in particular $m_{0011}=m_{0101}$. Imposing this condition on $P$ reduces it to a polynomial with $2060$ terms. If we further assume a symmetric i.i.d. distribution and define $\mu_{4}=m_{0000}=\mathbb{E}[x^{4}]$ and $\mu_{2}^{2}=m_{0011}=m_{0111}=\mathbb{E}[x^{2}]^{2}$, the discriminant specializes to a polynomial beginning as
$729\mu_{2}^{24}t_{00}^{3}t_{11}^{3}+1458\mu_{2}^{22}\mu_{4}t_{00}^{4}t_{11}^{2}+1458\mu_{2}^{22}\mu_{4}t_{00}^{2}t_{11}^{4}-2187\mu_{2}^{20}\mu_{4}^{2}t_{00}^{4}t_{01}^{2}+\cdots$
The full expression is given in the appendix. Similar to the previous example, slices of the teacher-metric data discriminant can be visualized as varying focal loci (Figure 2, right).
@@ 4 Networks with Quadratic Activations
Section overview. We specialize our analysis to networks with quadratic activations, where the function space corresponds to symmetric matrices. This setting allows for sharper results on critical points and loss landscapes. We describe how parameter and function space critical points relate, characterize teacher-student losses as low-rank matrix approximation problems under distribution-induced norms, and present variations of Eckart-Young theorem for Frobenius and Gaussian norms, proved using geometric insights from the previous section. Finally, we show that non-Gaussian norms can lead to exponentially many critical points, highlighting the landscape’s strong dependence on the data distribution.
For $\mathcal{W}=(\alpha,W)\in\mathbb{R}^{r}\times\mathbb{R}^{r\times n}$, we write
@@ 4.1 Critical points of smooth loss functions
We begin by specializing and refining some of the general results on the critical points of smooth loss functions from Section 2 to the case of $d=2$. In the following, we write $\mathcal{S}(r;n)$ for the set of $n\times n$ symmetric matrices with rank exactly $r$. The set $\mathcal{S}(r;n)$ is a smooth manifold of dimension $\frac{1}{2}r(2n-r+1)$ and the embedded tangent space at a point $S\in\mathcal{S}(r;n)$ is $T_{S}\mathcal{S}(r;n)=\{SX+X^{\top}S\colon X\in\mathbb{R}^{n\times n}\}$ [helmke1995].
@@
We show that $T_{S}\mathcal{S}({r;n})^{\perp}=\{A\in S^{2}(\mathbb{R}^{{n}})\colon A\cdot S=0\}$. Indeed, if $A\cdot S=0$, then from the cycle property of the trace operator $\langle A,SX+{X^{T}}S\rangle_{F}=tr(ASX)+tr(AX^{\top}S)=0$ for any $X\in\mathbb{R}^{d\times d}$. Conversely, if $\langle A,SX+X^{\top}S\rangle_{F}=0$ for all $X$, then $As_{i}=0$ for all columns $s_{i}$ of $S$ which means that $A\cdot S=0$. ∎
@@
If $\ell:{\rm Sym}^{2}(\mathbb{R}^{n})\rightarrow\mathbb{R}$ is any smooth function and $L=\ell\circ\tau_{r}$, then for any $\mathcal{W}=(\alpha,W)\in\mathbb{R}^{r}\times\mathbb{R}^{r\times n}$ we have that
@@
If $L=\ell\circ\tau_{r}$ and $\mathcal{W}=(\alpha,W)$ is a critical point for $L$ such that $S=\tau_{r}(\mathcal{W})$, then $\nabla\ell(S)\cdot S=0$ holds. In particular, if $r^{\prime}=rk(S)$, we have that $S$ is a critical point of the restriction of $\ell$ to $\mathcal{S}(r{{}^{\prime}};n)$. More generally, the image of the entire critical set $\tau_{r}(\mathrm{Crit}(L))$ is equal to $\bigcup_{r^{\prime}\leq r}\mathrm{Crit}(\ell|_{\mathcal{S}(r^{\prime};n)})$.
@@
The first statement is a consequence of the first part of Proposition 31, noting that if $\mathcal{W}$ is a critical point for $L$ then $0=W\cdot\nabla_{W}L(\mathcal{W})=4\cdot S\cdot\nabla{\ell}(S)$. The second statement then follows from Proposition 30. For the last statement, consider $S$ such that $rk(S)=r^{\prime}<r$ and $\nabla\ell(S)\cdot S=0$, so $S\in\mathrm{Crit}(\ell|_{\mathcal{S}(r^{\prime};n)})$. A parameterization $S=\tau_{r}(\mathcal{W})$ such that $\alpha_{i}\neq 0$ for exactly $r^{\prime}$ indices will be such that then $\nabla L(\mathcal{W})=0$ (again from Proposition 31). It is enough to note that any $S$ with $rk(S)=r^{\prime}<r$ admits a representation in parameter space with exactly $r^{\prime}$ non-zero neurons. ∎
@@ 4.2 Teacher-student problems
The expression (28) essentially appears in [gamarnik2020, Theorem 3.1(b)], but as a norm rather than as an inner product. Using (27) and the assumptions on $\mathcal{D}$, we have that
@@
From Corollary 32, it is sufficient to describe critical points of the loss restricted to the low-rank manifold $\mathcal{S}(r;n)$. Thus, we consider teacher-student problems in function space:
@@
where $T\in{\rm Sym}^{2}(\mathbb{R}^{n})$ is fixed and $\|\cdot\|$ is a norm induced by one of the inner products above.
For the Frobenius inner product, the optimization landscape is described by the following version of Eckart-Young’s Theorem, shown in [helmke1995]. We present here a much shorter geometric proof using the general description of critical points for distance functions given in Section 3 (the proof in [helmke1995] consists in more than 3 pages of computations). We make use of the following intuitive description of the focal locus.
@@
Let $X=\mathcal{S}(r,n)$ and let $\Sigma$ correspond to the Frobenius or Gaussian norm. Given $S\in X$, consider $A\in{\rm Sym}^{2}(\mathbb{R}^{n})$ such that $\{T_{\gamma}=S+\gamma A\colon\gamma\geq 0\}$ is a $\Sigma$-orthogonal ray to $X$. Assume further that there exists a matrix along this ray with distinct eigenvalues. Then $T_{\gamma}$ is a focal point for $(X,S)$ if and only if it has a repeated eigenvalue. More precisely, the multiplicity of $T_{\gamma}=\sum_{i=1}^{n}\sigma_{i}(u_{i}\otimes u_{i})$ as a focal point is given by $\{(j,k)\colon j<k,\sigma_{j}=\sigma_{k}\}$.
@@
  • The function $h_{T}|_{\mathcal{S}(r;n)}:\mathcal{S}(r;n)\rightarrow\mathbb{R}$ has exactly $\binom{n}{r}$ critical points, given by $S_{\mathcal{I}}=U\Sigma_{\mathcal{I}}U^{T}$ for any $\mathcal{I}\subset\{1,\ldots,n\}$ such that $\mathcal{I}=r$.
@@
We have that $S\in{\mathcal{S}(r;n)}$ is a critical point of $h_{T}|_{\mathcal{S}(r;n)}$ if and only if $\langle T-S,XS+SX\rangle=0$ for all $X\in{\rm Sym}^{2}(\mathbb{R}^{n})$, which in turn is equivalent to $S\cdot(T-S)=0$ by Proposition 30. Now let $S=\sum_{i=1}^{r}\sigma_{i}^{\prime}(u_{i}^{\prime}\otimes u_{i}^{\prime})$ and $(T-S)=\sum_{j=1}^{s}\sigma_{j}^{\prime\prime}(u_{j}^{\prime\prime}\otimes u_{j}^{\prime\prime})$ be eigendecompositions of $S$ and TST-S with $\sigma_{i}^{\prime}\neq 0$ and $\sigma_{j}^{\prime\prime}\neq 0$. The row/column spaces of $S$ and TST-S are spanned by the vectors $u_{i}^{\prime}$ and $u_{j}^{\prime\prime}$, respectively. From the condition S(TS)=0S(T-S)=0, we deduce that $S$ is a critical point if and only if $u_{i}^{\prime}\cdot u_{j}^{\prime\prime}=0$ for all $i,j$ in the two decompositions, which means that
@@
is an eigendecomposition of $T$. This shows that critical points are of the form $S=U\Sigma_{\mathcal{I}}U^{T}$ where $U\Sigma U^{T}$ is an eigendecomposition of $T$ and $\mathcal{I}\subset\{1,\ldots,n\}$ is such that $\mathcal{I}=r$.
@@
  • The function $h_{T}|_{\mathcal{S}(r;n)}:\mathcal{S}(r;n)\rightarrow\mathbb{R}$ has exactly $\binom{n}{r}$ critical points, given by $S_{\mathcal{I}}=U\Lambda_{\mathcal{I}}U^{T}$ for any $\mathcal{I}\subset\{1,\ldots,n\}$ such that $\mathcal{I}=r$, with
@@
Assume that $S=\sum_{i=1}^{r}\sigma_{i}^{\prime}(u_{i}^{\prime}\otimes u_{i}^{\prime})$ and $(S-T)=\sum_{i=1}^{n}\sigma_{i}^{\prime\prime}(u_{i}^{\prime\prime}\otimes u_{i}^{\prime\prime})$. Setting $\tau=tr(S-T)$, we have $\nabla h_{T}(S)=\sum_{i=1}^{n}(2\sigma_{i}^{\prime\prime}+\tau)(u_{i}^{\prime\prime}\otimes u_{i}^{\prime\prime})$. From the condition $S\cdot\nabla h_{T}=0$, we deduce that $S$ and $\nabla h_{T}$ have orthogonal eigenvectors, i.e., $u_{i}^{\prime}\cdot u_{j}^{\prime\prime}=0$ holds for all $i$ and all $j$ such that $2\sigma^{\prime\prime}_{j}+\tau\neq 0$. This implies that a decomposition of (ST)(S-T) is such that, up to permutation of indices, $u_{i}^{\prime\prime}=u_{i}^{\prime}$ and $\sigma_{i}^{\prime\prime}=-\tau/2$ holds for $i=1,\ldots,r$, while $\sum_{i=r+1}^{n}\sigma^{\prime\prime}_{i}=\tau(r+2)/2$ (so that $tr(S-T)=\tau$). We thus have that an eigendecomposition of $T$ is given by
@@
To prove the second point, we use again Theorem 27 and Proposition 34. The eigenvalues of $M_{\alpha}=\alpha T+(1-\alpha)S$ are the diagonal elements of $\Sigma_{\mathcal{I}}+\alpha\Sigma_{\mathcal{I}^{c}}+(1-\alpha)c_{\mathcal{I}}I_{\mathcal{I}}$ where $c_{\mathcal{I}}=\frac{tr(\Sigma_{\mathcal{I}^{c}})}{r+2}$ and $I_{\mathcal{I}}$ is the diagonal matrix with $I_{\mathcal{I},ii}=1$ if $i\in\mathcal{I}$ and $I_{\mathcal{I},ii}=0$ otherwise. Two eigenvalues of $M_{\alpha}$ are the same if and if only there exist $i\in\mathcal{I}$ and $j\in\mathcal{I}^{c}$ such that $\sigma_{i}+(1-\alpha)c_{\mathcal{I}}=\alpha\sigma_{j}$, or equivalently $\alpha=\frac{\alpha_{i}+c_{\mathcal{I}}}{\alpha_{j}+c_{\mathcal{I}}}$. The number of intersections with the discriminant locus is thus equal to the number of pairs $(\sigma_{i},\sigma_{j})\in\mathcal{I}\times\mathcal{I}^{c}$ such that ${(\sigma_{i}+c_{\mathcal{I}})}/{(\sigma_{j}+c_{\mathcal{I}})}\in[0,1]$. ∎
A statement analogous to Theorem 36 can be derived in the same way for any matrix inner product of the form $\langle S,T\rangle=\alpha{tr}(ST)+\beta{tr}(S)\mathrm{tr}(T)$, because Proposition 34 also applies in these cases. In particular, this yields a complete description of the landscape for all distribution-based inner products associated with distributions satisfying $\mu_{4}=3\mu_{2}^{2}$. However, for more general inner products, there may be exponentially more critical points.
@@
has $(3^{n}-1)/2$ critical points whenever $T=\operatorname{diag}(t)$ for $t\in\mathcal{T}$.
Theorem 37 is proven in the appendix. This stands in contrast to the Gaussian and Frobenius cases for low-rank approximation, where the corresponding optimization problem has only $n$ critical points generically. The number $(3^{n}-1)/2$ in Theorem 37 coincides with the “generic ED degree” reported in [draisma2013, Example 5.6]. A numerical experiment illustrating these results is presented in the appendix.